How much will Avagadro's number of chlorine atoms weigh
1
Expert's answer
2012-02-23T09:42:30-0500
Divide the number of atoms(6.022*1023) by avogadros number (6.022*1023), the resulting number(1) is the number of moles you have. Multiply the number of moles of atoms(1) by the molar mass of chlorine(35.5) and the answer is how many grams(35.5g) of the substance you have.
n(Cl)=6.022*1023/6.022*1023=1 mol
m(Cl)=n(Cl)*M(Cl)=1mol*35.5g/mol=35.5g